OpenStudy (anonymous):

2,10,30,68,130,.......... what is the next number??

OpenStudy (anonymous):

222

OpenStudy (anonymous):

brilliant

OpenStudy (anonymous):

thanks

OpenStudy (lalaly):

(n^3 + n ) = 6^3+6=222